### v1

# FreeChunker: A Cross-Granularity Chunking Framework

Abstract
Chunking strategies significantly impact the effectiveness of Retrieval-Augmented Generation (RAG) systems. Existing methods operate within fixed-granularity paradigms that rely on static boundary identification, limiting their adaptability to diverse query requirements. This paper presents FreeChunker, a Cross-Granularity Encoding Framework that fundamentally transforms the traditional chunking paradigm: the framework treats sentences as atomic units and shifts from static chunk segmentation to flexible retrieval supporting arbitrary sentence combinations. This paradigm shift not only significantly avoids the computational overhead required for semantic boundary detection, but also enhances adaptability to complex queries. Experimental evaluation on LongBench V2 demonstrates that FreeChunker achieves superior retrieval performance compared to traditional chunking methods, while significantly outperforming existing methods in computational efficiency.
Machine Learning, Natural Language Processing, Information Retrieval, Chunking
1 Introduction
The rapid advancement of large language models (LLMs) has substantially propelled progress in natural language processing, enabling strong generalization capabilities across diverse tasks (Brown et al., 2020). However, LLMs typically rely on static, parameterized knowledge, often making them less suitable for scenarios requiring up-to-date or domain-specific information. This limitation frequently leads to hallucinated content that may deviate from factual correctness (Huang et al., 2025; Li et al., 2023).
To address this limitation, RAG has emerged as a promising paradigm. By incorporating external knowledge sources, RAG enhances the generation process with more timely and verifiable content, achieving strong performance on knowledge-intensive tasks such as open-domain question answering and fact verification (Lewis et al., 2020). However, the effectiveness of RAG critically depends on the quality of retrieved content, which is substantially influenced by document chunking strategies (Ru et al., 2024). Prior research has demonstrated that chunking granularity often directly affects intra-chunk coherence, semantic coverage, and downstream generation quality. Inappropriate chunking strategies may disrupt semantic structures, potentially leading to information fragmentation or redundancy, ultimately undermining both retrieval and generation performance (Brådland et al., 2025; Lyu et al., 2025).
Figure 1: Cross-Granularity Chunking FreeChunker breaks the limitation of fixed-granularity text chunking, enabling flexible extraction of sentence combinations at arbitrary granularities.
To improve chunking quality, recent efforts have explored semantic-aware chunking techniques. For instance, SemanticChunker (LangChain, 2024) segments documents based on sentence embedding similarity. Meta-Chunking (Zhao et al., 2024) identifies semantic boundaries through perplexity analysis and probability-based segmentation decisions. LumberChunker (Duarte et al., 2024) leverages LLMs to predict chunk boundaries more aligned with human interpretation. However, existing approaches often lack the flexibility to adapt to diverse query intents or varying deployment constraints. Moreover, some methods rely on computationally expensive large models, where the marginal gains of semantic chunking may not always justify the computational overhead (Qu et al., 2025). More fundamentally, these approaches still operate within a single-granularity paradigm, consequently limiting their ability to simultaneously accommodate different information needs within the same document collection.
To overcome the fundamental limitation of fixed chunking schemes that cannot simultaneously provide multiple granularity levels, FreeChunker is proposed in this paper—a novel Cross-Granularity Encoding Framework that introduces an alternative chunking paradigm. As illustrated in Figure 1, unlike traditional chunking methods that are typically constrained by fixed text boundaries, the proposed method enables retrievers to extract arbitrary combinations of sentences from documents, thereby providing enhanced flexibility in content retrieval. The following core contributions are offered by the proposed method:
  • Cross-granularity chunking paradigm: A novel approach is proposed that breaks the limitation of fixed-granularity text chunking, enabling flexible extraction of sentence combinations at arbitrary granularities.
  • Computational efficiency: The framework reduces chunking overhead by treating sentences as atomic units, avoiding expensive boundary identification processes.
  • Superior performance: The method achieves competitive retrieval performance by adapting multiple granularity levels to diverse query requirements, demonstrating enhanced effectiveness across various information retrieval scenarios in LongBench V2 Task.
Figure 2: The architecture of FreeChunker. Long text input is processed by the Sentenizer and Embedding Model to produce sentence embeddings. The Attention Matrix and Chunk Pattern Mask work together to generate multiple granularity chunk embeddings in a single forward pass, enabling flexible on-demand access to different chunk sizes.
2 Related Work
Limitations of Fixed-granularity Chunking
As a foundational step in RAG, text chunking substantially influences overall system performance. The most straightforward strategy divides documents into fixed-granularity chunks, which is simple to implement but often misaligns with semantic boundaries, consequently leading to information fragmentation or redundancy (Ru et al., 2024). More critically, fixed chunking forces a single granularity choice across the entire document, thereby preventing the system from simultaneously accessing both fine-grained details and coarse-grained context that different queries may require. This granularity rigidity fundamentally limits the retrieval system’s ability to adapt to diverse information needs within a single document collection.
Semantic-aware Chunking Techniques
To address the aforementioned semantic boundary issues, recent work has explored more adaptive chunking strategies. SemanticChunker (LangChain, 2024) relies on sentence-level embedding similarity to enhance intra-chunk semantic coherence by identifying natural breakpoints based on semantic similarity thresholds. Meta-Chunking (Zhao et al., 2024) proposes two adaptive segmentation algorithms leveraging large language models’ logical awareness: Perplexity Chunking identifies logical boundaries by analyzing inter-sentence perplexity variations, while Margin Sampling Chunking determines boundaries by comparing model probability differences for segmentation decisions. LumberChunker (Duarte et al., 2024) leverages large language models to determine chunk boundaries more aligned with human interpretation. However, despite their semantic awareness, these methods still operate under a single granularity paradigm—each document is ultimately represented at one chosen level of granularity, whether sentence-level, paragraph-level, or custom semantic units.
Adaptive Attempts
Inspired by the concept of mixture of experts (MoE) (Jacobs et al., 1991; Jordan & Jacobs, 1994), recent methods like MoC (Zhao et al., 2025) and MoG (Mix-of-Granularity) (Zhong et al., 2025) attempt to enhance adaptability by treating different chunking strategies as expert modules and employing routing mechanisms to dynamically select the most suitable chunker based on query characteristics. However, these MoE-inspired methods essentially aggregate existing chunkers as experts; consequently, both the quality of semantic chunking and the degree of granularity control remain fundamentally bounded by the capabilities of the underlying chunker experts.
Despite these advances, a fundamental limitation persists: current methods still operate within a single-chunking, single-granularity paradigm—whether fixed, semantics-driven, or query-adaptive. Consequently, retrieval systems cannot simultaneously access multiple levels of information granularity, which is precisely what is needed to better accommodate diverse query intents. Therefore, there remains a critical need for a paradigm that can transcend these single-granularity constraints.
3 Methodology
In FreeChunker, sentences are treated as atomic units and chunk embeddings are generated across multiple granularities simultaneously. This approach typically enables retrieval systems to access any desired chunk size on demand, ranging from fine-grained single sentences to coarse-grained multi-sentence contexts.
3.1 Cross-Granularity Chunk Pattern
The core innovation of FreeChunker lies in proposing a flexible multi-granularity paradigm that overcomes the fixed granularity limitation of existing methods by defining multiple chunk patterns, thereby enabling on-demand granularity access.
Given a document with $n$ sentences $\mathcal{D}=\{s_{1},s_{2},\dots,s_{n}\}$, a Chunk Pattern Mask $\mathbf{P}\in\{0,-\infty\}^{m\times n}$ is constructed, where each row $\mathbf{P}[i,:]$ represents the $i$-th chunk pattern, and $\mathbf{P}[i,j]=0$ typically indicates that sentence $s_{j}$ belongs to that pattern.
The key insight is that the matrix can encode arbitrary granularity combinations. For any granularity $g$ and starting position $s$, chunk patterns are defined as:
$\mathbf{P}_{g,s}[i,j]=\begin{cases}0&\text{if }s\leq j<s+g\\ -\infty&\text{otherwise}\end{cases}$ (1)
Therefore, the framework can support any combination of chunk sizes and overlapping strategies.
3.2 Cross-Granularity Encoder
After obtaining the Chunk Pattern Mask $\mathbf{P}$, independently encoding each chunk would typically cause extensive redundant computation. Sentenizer is introduced to split documents into sentences and encode them using a token-level embedding model $\mathcal{M}$, thereby obtaining the sentence embedding matrix $\mathcal{E}=[\vec{e_{1}},\dots,\vec{e_{n}}]\in\mathbb{R}^{n\times d}$. For any chunk spanning from sentence $s_{i}$ to sentence $s_{j}$, its true embedding is denoted as $\vec{e}_{i,j}=\mathcal{M}([s_{i},s_{i+1},\dots,s_{j}])$, where $s_{i},s_{i+1},\dots,s_{j}]$ represents the concatenation of sentences.
A specialized transformer architecture is then modified. For each layer, a learnable chunk embedding $\mathbf{h}_{\text{chk}}\in\mathbb{R}^{d}$ is introduced that is replicated $m$ times to form $\mathcal{H}=[\mathbf{h}_{\text{chk}},\mathbf{h}_{\text{chk}},\dots,\mathbf{h}_{\text{chk}}]^{T}\in\mathbb{R}^{m\times d}$ and attention is used to generate $m$ chunk embeddings in parallel:
$\mathbf{Q}=\mathbf{W}_{Q}\mathcal{H},\quad\mathbf{K}=\mathbf{W}_{K}\mathcal{E},\quad\mathbf{V}=\mathbf{W}_{V}\mathcal{E}$ (2)
$\text{Attn}(\mathcal{H},\mathcal{E})=\mathrm{softmax}\!\Big(\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}+\mathbf{P}\Big)\mathbf{V}$ (3)
The complete chunk encoding layer then applies layer normalization and feed-forward networks:
$\mathcal{V}=\text{LayerNorm}(\mathcal{H}+\text{Attn}(\mathcal{H},\mathcal{E}))$ (4)
$\text{Output}=\text{LayerNorm}(\mathcal{V}+\text{FFN}(\mathcal{V}))$ (5)
This design typically allows parallel generation of all chunk embeddings in a single forward pass, with sentence-level embeddings being reused across different granularities, thereby avoiding redundant encoding of sentences.
Figure 3: Training Dynamics Across Model Scales. Training and validation curves (top) and test set loss distributions (bottom) for three different model configurations, demonstrating consistent convergence and effective generalization.
4 Theoretical Analysis
A central concern for FreeChunker is whether the cross-granularity encoder induces non-negligible substitution loss relative to directly encoding concatenated text chunks. This section develops a geometric analysis of vector substitution, establishing worst-case and average-case guarantees that are verifiable in practice.
4.1 Setup and Constraints
Let $\vec{e}_{i,j}=\mathcal{M}([s_{i},s_{j}])$ denote the true chunk embeddings of the concatenated chunk $s_{i},s_{j}]$, and let $\vec{v}_{i,j}$ be the approximate chunk embeddings produced by the cross-granularity encoder. Let $\vec{q}$ be a query embedding. Throughout this section all vectors are assumed to be unit-norm.
The substitution loss is studied as follows:
$\varepsilon\;:=\;\big|\cos(\vec{q},\vec{e}_{i,j})-\cos(\vec{q},\vec{v}_{i,j})\big|.$ (6)
Two parameters characterize the geometric configuration:
$\rho$ $\;=\;\cos(\vec{e}_{i,j},\vec{v}_{i,j}),$ (7)
$s$ $\;=\;\cos(\vec{q},\vec{e}_{i,j}),$
where $\rho$ quantifies substitution quality and $s$ captures query alignment with the true chunk.
4.2 Worst-Case Substitution Bound
Theorem 1 (Vector Substitution Loss Upper Bound). Given the parameters in (7),
$\varepsilon\;\leq\;s(1-\rho)\;+\;\sqrt{1-s^{2}}\,\sqrt{1-\rho^{2}}.$ (8)
Proof. Let $\alpha=\angle(\vec{q},\vec{e}_{i,j})$ and $\beta=\angle(\vec{e}_{i,j},\vec{v}_{i,j})$, so $s=\cos\alpha$ and $\rho=\cos\beta$. By the spherical cosine law,
$\cos(\vec{q},\vec{v}_{i,j})\;=\;\cos\alpha\cos\beta+\sin\alpha\sin\beta\cos\varphi,$ (9)
where $\varphi$ is the relative azimuth of $\vec{q}$ and $\vec{v}_{i,j}$ around the axis $\vec{e}_{i,j}$. The adversarial configuration is $\cos\varphi=-1$, which yields $\cos(\vec{q},\vec{v}_{i,j})\geq s\rho-\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}$. Subtracting from $\cos(\vec{q},\vec{e}_{i,j})=s$ gives (8). $\square$
Interpretation.
The first term $s(1-\rho)$ is typically the radial component; the second term $\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}$ is tangential and becomes dominant when the query is weakly aligned with the true chunk ($s$ small) and the substitution deviates from the truth ($\rho$ small).
4.3 Expected Error under Random Azimuth
The worst-case bound (8) assumes an adversarial azimuth between $\vec{q}$ and $\vec{v}_{i,j}$. In practice, the relative azimuth $\varphi$ between these vectors around $\vec{e}_{i,j}$ is typically random rather than adversarial. A mild and empirically verifiable assumption is therefore introduced.
Assumption 1 (Random Azimuth). The relative orientation between query and approximate vectors around the true chunk embeddings is uniformly random.
Theorem 2 (Expected Substitution Loss). Under Assumption 1 and given the parameters in (7), the expected substitution loss is bounded by:
$\mathbb{E}_{\varphi}[\varepsilon\mid s,\rho]\leq s(1-\rho)+\frac{2}{\pi}\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}.$ (10)
Proof. Using $\varepsilon\leq s(1-\rho)+\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}|\cos\varphi|$ and $\mathbb{E}[|\cos\varphi|]=2/\pi$ for uniform $\varphi$ yields the bound. $\square$
Interpretation.
The bound consists of a radial term $s(1-\rho)$ and a tangential term with coefficient $2/\pi\!\approx\!0.636$, showing that random orientations substantially reduce expected degradation compared to worst-case scenarios.
Table 1: Results of different chunking methods on LongBench V2 (without Top-1 results)
Method jina-embeddings-v2-small-en nomic-embed-text-v1.5 BGE-M3
Top-5 Top-10 Time (↓) Top-5 Top-10 Time (↓) Top-5 Top-10 Time (↓)
I. Single-Document QA
Traditional 30.86% 30.86% 2.11s 35.43% 32.00% 3.76s 36.00% 34.29% 8.93s
SemanticChunker 33.71% 33.71% 12.87s 29.14% 32.57% 15.48s 32.57% 30.86% 19.83s
PPL Chunking 31.43% 28.00% 33.09s 32.57% 34.29% 34.73s 33.14% 33.71% 39.82s
Margin Sampling 30.29% 36.57% 364.64s 35.43% 36.00% 368.68s 34.29% 28.00% 375.22s
FreeChunker 32.57% 36.57% 2.75s 36.00% 36.00% 5.29s 38.29% 36.57% 9.64s
5 Training Process
5.1 Training Setup
Token-level Embedding Models
Three token-level embedding models are employed as encoders $\mathcal{M}$ for the Sentenizer: jina-embeddings-v2-small-en(Günther et al., 2023)(33M parameters), nomic-embed-text-v1.5 (Nussbaum et al., 2024) (137M parameters) and BGE-M3 (Chen et al., 2024) (570M parameters). Our sentence-level Cross-Granularity encoder (330M parameters) reuses partial structures and weights from BGE-M3, with dimension pruning support enabling token-level embedding model reuse across different dimensions. Documents are split into sentences and encoded to generate sentence embeddings $\mathcal{E}=[\vec{e_{1}},\dots,\vec{e_{n}}]\in\mathbb{R}^{n\times d}$.
Dataset and Preprocessing
The training dataset comprises 100K documents sampled from The Pile (Gao et al., 2020). For each document, the first 8000 tokens are extracted to construct training pairs. Documents are split into sentences using the Sentenizer, typically yielding 10-200 sentences per document. Training pairs $(\vec{e},\vec{v})$ are constructed across 5 different granularities (2, 4, 8, 16, 32 sentences), where $\vec{e}$ represents ground-truth chunk embeddings obtained by directly encoding concatenated sentence spans, and $\vec{v}$ denotes the corresponding cross-granularity encoder outputs.
Optimization
The training objective is formulated to minimize the cosine similarity loss:
$\mathcal{L}=1-\frac{1}{|\mathcal{B}|}\sum_{(\vec{e},\vec{v})\in\mathcal{B}}\cos(\vec{e},\vec{v})$ (11)
where $\mathcal{B}$ denotes the training batch. The AdamW optimizer is employed with learning rate $1\times 10^{-4}$ and batch size 1. Training is conducted for 2 epochs with cosine schedule learning rate decay, incorporating warmup over the first one-third of total training steps. Model validation is performed every 1000 steps on a held-out validation set of 200 samples to monitor generalization performance.
Computational Resources
All training experiments are conducted on NVIDIA A800 GPUs with 80GB memory. The total computational cost exceeds 500 GPU hours, encompassing dataset encoding construction and model training processes.
5.2 Training Dynamics and Evaluation
Convergence Analysis
The upper panels of Figure 3 present the training and validation cosine loss curves for three base models. All configurations demonstrate smooth and stable convergence within 200K steps, validating the effectiveness of the overall optimization strategy. Notably, the smallest model, jina-embeddings-v2-small-en (33M parameters), exhibits the fastest and cleanest convergence with lower training loss. As model parameters increase, both training and validation losses show degradation across the three models. This phenomenon may be attributed to the use of a fixed 330M-parameter sentence-level encoder for all three base models, suggesting that scaling the sentence-level encoder capacity could potentially improve fitting performance for BGE-M3 (570M parameters).
Test Set Generalization Analysis
The lower panels of Figure 3 illustrate the cosine loss distributions on the independent test set. We employ a threshold of 0.06 to distinguish well-fitted samples from under-fitted ones, enabling coarse-grained assessment of chunk embedding accuracy. For jina-embeddings-v2-small-en, 100% of test samples fall below this threshold, with loss distribution highly concentrated in the 0.01–0.02 range, reflecting excellent generalization capability and minimal approximation error. In contrast, nomic-embed-text-v1.5 (137M parameters) exhibits a more dispersed distribution: 96.7% of samples remain below the 0.06 threshold, yet long-tail high-loss samples emerge, indicating increased representation noise or model uncertainty. BGE-M3 (570M parameters) shows further distribution broadening, consistent with the training loss patterns observed.
6 Experiments
6.1 Experimental Setup
Datasets
Comprehensive experiments are conducted on LongBench V2 (Bai et al., 2024), a challenging benchmark specifically designed for long-context understanding tasks. This dataset is suitable for evaluating chunking methods as it contains documents with extensive contexts and covers diverse task types with varying difficulty levels, thereby providing a realistic testbed for assessing the effectiveness of different chunking strategies in practical RAG scenarios.
Embedding Models
To ensure fair comparison, identical text encoding models are utilized across all methods. Three embedding models are employed for comprehensive evaluation: jina-embeddings-v2-small-en (33M parameters), nomic-embed-text-v1.5 (137M parameters), and BGE-M3 (570M parameters). These models typically represent different scales and architectures, thereby enabling thorough assessment of chunking performance across varying embedding capacities.
Generative Model
For text generation, Qwen3-8B (Team, 2025) is employed as the backbone language model, deployed via vLLM for efficient inference. To guarantee fair comparison and reproducibility, greedy decoding with temperature set to 0 is employed by Qwen3-8B.
Baseline Methods
FreeChunker is compared against several representative chunking approaches. (1) Traditional Chunking represents the traditional approach that first splits text by periods to preserve semantic integrity, then iteratively accumulates sentences until exceeding predefined token limits (256 tokens); (2) Semantic Chunking employs SemanticChunker (LangChain, 2024) to perform embedding similarity-based chunking using sentence embeddings to identify natural breakpoints in the text. The breakpoint threshold type set to "percentile" and breakpoint threshold amount set to 50.0; (3) Meta-Chunking (Zhao et al., 2024) employs perplexity-based semantic boundary detection with two sub-methods: PPL Chunking and Margin Sampling, utilizing Qwen/Qwen2.5-1.5B-Instruct model (Yang et al., 2024) for text chunking. PPL Chunking adopts the threshold of 0.5 as recommended in the original paper, while Margin Sampling uses default settings; (4) LumberChunker (Duarte et al., 2024) provides LLM-guided chunk boundary prediction; however, due to excessive processing time for long texts, it is excluded from comparisons involving lengthy documents.
6.2 Main Results
Table 1 presents the results across three embedding models on LongBench V2. FreeChunker demonstrates superior performance in complex query environments while achieving significant computational efficiency gains compared to traditional approaches.
Regarding computational efficiency, FreeChunker maintains processing time comparable to traditional chunking methods while delivering superior retrieval performance. In contrast, existing semantic chunking approaches require significantly longer processing time without achieving comparable performance gains. This combination of enhanced effectiveness and computational efficiency validates the core contributions of the cross-granularity paradigm in practical deployment scenarios.
7 Conclusion
This paper introduces FreeChunker, a novel cross-granularity chunking framework that fundamentally transforms the traditional text chunking paradigm. By shifting from static boundary identification to flexible semantic combination retrieval, the proposed method significantly reduces computational overhead associated with semantic analysis for boundary detection while providing enhanced adaptability to diverse query requirements. Unlike conventional static chunking approaches, this flexible semantic combination strategy accommodates varying and unpredictable chunking demands across different query intents.
Beyond the adjacent sentence combination approach explored in our experiments, the framework’s chunk pattern control mechanism $\mathbf{P}_{g,s}[i,j]$ enables non-contiguous sentence linking capabilities. This opens promising research directions for incorporating predefined network structures within document collections to construct sophisticated non-adjacent sentence connections. Such extensions represent compelling avenues for future investigation, potentially enabling more complex semantic relationship modeling across document structures.
The experimental results demonstrate that FreeChunker achieves competitive retrieval performance while maintaining computational efficiency comparable to baseline methods. The framework’s ability to provide multiple granularity levels simultaneously addresses fundamental limitations of existing single-granularity approaches, offering enhanced flexibility for diverse information retrieval scenarios in practical RAG applications.